OpenStudy (anonymous):

Determine whether the given linear equations are parallel, perpendicular, or neither. y = -9x +11

OpenStudy (anonymous):

y=1/9x+8 and y = -9x +11

OpenStudy (anonymous):

Can u find slopes of the lines?

OpenStudy (anonymous):

yes 1/9x and -9x

OpenStudy (anonymous):

when you multiply them you get -1

OpenStudy (anonymous):

hence they are perpendicular

OpenStudy (anonymous):

oh thanks!

OpenStudy (anonymous):

Determine whether the given linear equations are parallel, perpendicular, or neither. 8x - 4y = - 12 8x - 4y = 8

OpenStudy (anonymous):

parallel right?

OpenStudy (anonymous):

if m1 is the slope of the 1st line and m2 is slope of second line if m1 = m2 then two lines are said to be parallel if m1*m2 = -1 then two are said to be perpendicular if none of above satisfies then two lines are neither parallel nor perpendicular

OpenStudy (anonymous):

Thank you! Write the equation of the line with a slope of 5 and a y-intercept of (0, 6) in slope-intercept form.